# A Note on Restricted Selection Set from Random Interval

### v1

@@ Abstract.
We study restricted selection sets of random intervals in $\mathbb{R}^{1}$ defined on a non-atomic probability space. Given a random interval $Y=[y_{L},y_{U}]$ and scalar constraints on the expectation or the median of admissible selections, we define the restricted selection set and establish its existence, basic structure and influence on bounding moments and quantiles. In particular, we give conditions under which any mean (or quantile) in the Aumann expectation range can be attained by a measurable selection. We characterize the induced ranges of means, medians, and event probabilities. The analysis is carried out in a minimal one-dimensional random-set framework inspired by the classical theory of Aumann integrals. We also outline extensions to higher-order moment and general quantile restrictions.
@@
Random sets, random intervals, Aumann expectation, selection, partial identification
@@ 1. Introduction
Let $(\Omega,\mathcal{A},\mathbf{P})$ be a complete non-atomic probability space and let $Y=[y_{L},y_{U}]$ be an interval-valued random closed set on $\mathbb{R}$, that is, a measurable map $Y:\Omega\to\mathcal{K}(\mathbb{R})$ taking values in compact intervals
$Y(\omega)=[y_{L}(\omega),y_{U}(\omega)],\qquad\omega\in\Omega,$
for some real-valued random variables $y_{L},y_{U}$ with $y_{L}\leq y_{U}$ almost surely. A measurable selection of $Y$ is a random variable $y:\Omega\to\mathbb{R}$ such that $y(\omega)\in Y(\omega)$ for $\mathbf{P}$–almost every $\omega$. The Aumann (selection) expectation of $Y$ is the closed set
@@
where $\operatorname{\mathbf{Sel}}^{1}(Y)$ denotes the family of integrable selections of $Y$; see [Aumann1965, Artstein1974, Molchanov2005]. In the one-dimensional interval case this set has the simple representation
$\mathbb{E}[Y]=\big[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})\big],$ (1.2)
so that the possible expectations of selections $y\in\operatorname{\mathbf{Sel}}(Y)$ form a compact interval determined solely by the expectations of the endpoints.
In this note, we ask how the bounds in (1.2) change once one restricts attention to a non-empty subset of selections $\mathcal{S}\subset\operatorname{\mathbf{Sel}}^{1}(Y)$. Specifically, we consider subsets of $\operatorname{\mathbf{Sel}}^{1}(Y)$ satisfying elementary scalar constraints. Concretely, we consider restricted selection sets defined by moment or quantile conditions, such as
$\operatorname{\mathbf{Sel}}(Y\mid\kappa):=\{y\in\operatorname{\mathbf{Sel}}(Y):\mathbb{E}[y]=\kappa\},\qquad\operatorname{\mathbf{Sel}}(Y\mid m):=\{y\in\operatorname{\mathbf{Sel}}(Y):Med(y)=m\},$
where $\operatorname*{Med}(y)$ denotes median of $y$. These sets of selections encode additional information about the law of the underlying latent variable $y^{*}$ constrained to lie in $Y(\omega)$ almost surely. Our aim is to understand (i) when such restricted selection sets are nonempty, (ii) show how restricting they can be in terms of the random interval $Y=[y_{L},y_{U}]$ and the underlying probability space, and (iii) which values of basic scalar functionals,
$\mathbb{E}[y],\qquad\operatorname*{Med}(y),\qquad\mathbf{P}(y\in A),$
can be realized by selections $y$ satisfying a given restriction.
The non-atomicity of $(\Omega,\mathcal{A},\mathbf{P})$ allows us to use measurable selection and rearrangement arguments in a sharp way. At a technical level, several of our proofs combine classical properties of the Aumann expectation with Hardy–Littlewood type inequalities and quantile representations. This yields explicit and, in many cases, closed-form descriptions of the admissible ranges of expectations, medians, event probabilities, and higher moments under simple one-dimensional constraints on the selections.
@@ 1.1. Related Literature
Our work is rooted in the classical theory of random closed sets and selection expectation. The notion of the Aumann integral for set-valued mappings was introduced in [Aumann1965] and developed further in [Artstein1974], where convexity and support-function representations play a central role. In the one-dimensional setting considered here, the random sets are intervals $Y=[y_{L},y_{U}]$, and the identity $\mathbb{E}[Y]=[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})]$ is a special case of these general results. For the basic theory of random closed sets, measurable selections, and the measurable graph theorem, we refer to [KuratowskiRyll1965, CastaingValadier1977, Molchanov2005].
Rearrangement and quantile methods appear throughout analysis and probability in the study of extremal inequalities. Our arguments for quantile attainability and median-restricted expectations are in the spirit of classical Hardy-Littlewood-Pólya inequalities, as presented for instance in [LiebLoss2001, HardyLittlewoodPolya1952]. We employ these tools in a simple random interval setting, where they yield explicit formulas for the effect of median constraints on the expectation and for extremal event probabilities under a mean restriction.
Random sets and Aumann expectations have been used in a variety of applied contexts, including mathematical economics, finance, and statistics; see, for example, the monograph [Molchanov2005] and the references therein. In particular, models with set-valued or interval-valued observations, or with latent variables constrained to lie in random sets, naturally lead to parameter ranges described in terms of measurable selections. While such applications motivate our interest in restricted selection sets, the present note is purely structural: we work throughout in a one-dimensional probability-theoretic setting and focus on existence, sharpness, and representation results for the restricted ranges of basic scalar functionals.
@@ 1.2. Our Contribution
This note develops a detailed description of restricted selection sets for random intervals on a non-atomic probability space. Starting from the benchmark Aumann expectation range and median interval for arbitrary selections $y\in\operatorname{\mathbf{Sel}}(Y)$, we first show that these bounds are sharp in a strong sense. For expectations, every point of $\mathbb{E}[Y]=[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})]$ is realized as $\mathbb{E}[y]$ for some measurable selection (Proposition 2.2). More generally, for each $\alpha\in(0,1)$, any value between the capacity and containment quantiles of $Y$ can be attained as the $\alpha$–quantile of a selection (Proposition 2.4).
We then introduce mean- and median-restricted selection sets and analyze the induced ranges of expectations. In the median-restricted case, we obtain sharp identification intervals for the mean: under mild feasibility conditions, the set
@@
We also examine the converse question of how mean restrictions affect the set of attainable medians. A simple two-state construction (Example 2.11) shows that a mean restriction $\mathbb{E}(y^{*})=\kappa$ does not, in general, reduce the benchmark median interval $\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})]$: every median in this interval can remain attainable under the mean constraint. For event probabilities $\mathbf{P}(y^{*}\in A)$, we refine the classical random-set bounds in Proposition 2.12 by imposing a mean restriction and deriving a dual (Lagrangian) representation (Theorem 2.13). This dual problem leads to an explicit threshold selection rule (Theorem 2.14) and a closed-form identification formula for $\mathbf{P}(y^{*}\in A)$ in terms of extremal distances between $Y$ and the target set $A$ (Corollary 2.15), while preserving consistency with the unrestricted capacity/containment bounds (Proposition 2.16).
Finally, we indicate extensions of the restricted selection framework in two directions. First, we consider higher-moment constraints and show how to obtain dual representations for the corresponding ranges of expectations using a simple two-dimensional random-set construction (Theorem 2.20). Second, we discuss quantile restrictions at levels $\alpha\neq\tfrac{1}{2}$ and describe the induced intervals of admissible means in this more general setting (Proposition 2.23). These extensions illustrate that the one-dimensional random-interval framework, combined with standard tools from the theory of random sets and rearrangements, already captures a rich collection of sharp identification and extremal results for restricted selections.
@@ 1.3. Notation and preliminaries
Spaces and sets. We denote by $\mathbb{R}$ the real line and, for $d\in\mathbb{N}$, by $\mathbb{R}^{d}$ the dd–dimensional Euclidean space equipped with its usual topology and Euclidean norm $\|\cdot\|$. Let $\mathcal{K}(\mathbb{R})$ be the family of all nonempty closed subsets of $\mathbb{R}$. Our random sets will always take values in $\mathcal{K}(\mathbb{R})$. For general background on random closed sets we refer to [Molchanov2005].
@@
where $\mathcal{B}(\mathcal{K}(\mathbb{R}))$ is the Borel $\sigma$–algebra generated by the Fell (hit–or–miss) topology (see [Molchanov2005]). In this note we consider the special case of random intervals, that is,
@@
The existence of measurable selections for closed-valued measurable multifunctions is classical, see e.g. [KuratowskiRyll1965, CastaingValadier1977]. We denote the family of all measurable selections of $Y$ by
@@
Following [Aumann1965], the Aumann expectation (selection expectation) of a random closed set $Y$ with at least one integrable selection is defined as the closed set
@@
where the closure is taken in the usual topology on $\mathbb{R}$; see also [Artstein1974, Molchanov2005]. For a random interval $Y=[y_{L},y_{U}]$ satisfying (1.3), one has the simple representation
@@
see [Molchanov2005]. In the one–dimensional setting we will mainly use the half–line sets $(-\infty,t]$, $t\in\mathbb{R}$, and write
@@
We write $\operatorname*{Med}(Z)$ for the (possibly set-valued) set of all such medians; in the scalar case $\operatorname*{Med}(Z)$ is always a nonempty closed interval (possibly degenerate).
@@
2. Main Result
In this section, we study restricted selection sets of a random interval $Y=[y_{L},y_{U}]$ on the non-atomic probability space $(\Omega,\mathcal{A},\mathbf{P})$ introduced in Subsection 1.3. We first record benchmark ranges for the expectation, median, and event probabilities of measurable selections of $Y$, and then investigate how these ranges are sharpened by simple scalar restrictions on the mean or median of the selection. Throughout we assume $y_{L},y_{U}$ are measurable real-valued random variables with $\mathbf{P}(y_{L}\leq y_{U})=1$ and
@@ 2. Main Result
2.1. Benchmark identification ranges
Consider a latent scalar random variable $y^{*}$ such that
$\mathbf{P}\big(y_{L}\leq y^{*}\leq y_{U}\big)=1.$
Equivalently, $y^{*}$ is a measurable selection of the random interval $Y=[y_{L},y_{U}]$. Without additional information on the distribution of $y^{*}$, the following benchmark identification ranges for its expectation and median are immediate.
@@ 2.1. Benchmark identification ranges
$\mathbb{E}(y^{*})\in\Theta_{E}^{*}:=\big[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})\big].$ (2.1)
Second, since $y_{L}\leq y^{*}\leq y_{U}$ almost surely, the median of $y^{*}$ must lie between the median of $y_{L}$ and the median of $y_{U}$. Writing $\operatorname*{Med}(Z)$ for the (possibly set-valued) set of medians of a real-valued random variable $Z$ (see Subsection 1.3), we obtain
$\operatorname*{Med}(y^{*})\in\Theta_{M}^{*}:=\big[\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})\big],$ (2.2)
where the right-hand side is understood as the smallest interval containing $\operatorname*{Med}(y_{L})$ and $\operatorname*{Med}(y_{U})$.
@@ 2.2. Restricted selection sets
Let $Y=[y_{L},y_{U}]$ be a random interval with measurable endpoints satisfying $y_{L}\leq y_{U}$ a.s. and assume $y_{L},y_{U}\in L^{1}$, the set of finitely integrable random variables. For each $t\in[0,1]$ define
@@ 2.3. Quantile attainability via selections
Let $(\Omega,\mathcal{F},\mathbf{P})$ be a non-atomic probability space and let $Y$ be a (measurable) random closed convex set in $\mathbb{R}$ (random interval). For $t\in\mathbb{R}$, define
@@
On $A^{-}\cup A^{0}$, the set $Y^{\leq m}(\omega)$ is nonempty; on $A^{0}\cup A^{+}$, the set $Y^{>m}(\omega)$ is nonempty. Because $Y$ has a measurable graph and the half-lines are Borel, $\operatorname{graph}(Y^{\leq m})$ and $\operatorname{graph}(Y^{>m})$ are measurable. By standard measurable selection results (e.g. via a Castaing representation, see [KuratowskiRyll1965, CastaingValadier1977]) we can take measurable selections
@@
with the additional requirement that $z^{-}(\omega)=\sup Y^{\leq m}(\omega)$ and $z^{+}(\omega)=\inf Y^{>m}(\omega)$ on their domains.
Since $Y(\omega)=[y_{L}(\omega),y_{U}(\omega)]$ is an interval, these extremal selections have a simple form:
$z^{-}(\omega)=\begin{cases}y_{U}(\omega),&\omega\in A^{-},\\[2.0pt] m,&\omega\in A^{0},\end{cases}\qquad z^{+}(\omega)=\begin{cases}m,&\omega\in A^{0},\\[2.0pt] y_{L}(\omega),&\omega\in A^{+}.\end{cases}$
In particular,
[…]
For notational convenience we complete these to global selections by setting
@@
Taking $\alpha=\tfrac{1}{2}$, the previous proposition implies that every point in $\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})]$ is a median of some selection of $Y$, so the benchmark median interval $\Theta_{M}^{*}$ is sharp.

### v2

# Bounds for Restricted Selections of Random Sets

@@ Abstract.
We study constrained selection sets of random closed sets defined on a non-atomic probability space. Given a random interval $Y=[y_{L},y_{U}]$ and scalar constraints on the expectation or the median of admissible selections, we characterize the restricted selection set and establish sharp bounds on the attainable ranges of means, medians, and event probabilities. In particular, we give conditions under which every value in the Aumann expectation range is realized as the mean of a measurable selection, and we obtain explicit formulas for the extremal expectations under median and higher-moment restrictions via rearrangement and convex-duality arguments. We further show that the selection set of any random compact convex set in $\mathbb{R}^{d}$ can be approximated in $L^{1}$ by selection sets of disjoint unions of random cubes, each of which decomposes coordinate-wise into one-dimensional interval selection problems. This gives us an approximation-based reduction of constrained selection problems for random compact convex sets in $\mathbb{R}^{d}$.
@@
Random convex sets, Aumann expectation, Measurable selection, Sharp bounds
2020 Mathematics Subject Classification:
Primary 28B20, 60D05; Secondary 54C60, 26D15, 49J53
@@ 1. Introduction
Let $(\Omega,\mathcal{A},\mathbf{P})$ be a complete non-atomic probability space and let $Y:\Omega\to\mathcal{K}_{c}(\mathbb{R}^{d})$ be a random compact convex set. A measurable selection of $Y$ is a random vector $y:\Omega\to\mathbb{R}^{d}$ such that $y(\omega)\in Y(\omega)$ for $\mathbf{P}$–a.s. The family of all measurable selections is denoted $\operatorname{\mathbf{Sel}}(Y)$, and the Aumann expectation of $Y$ is the closed set
@@
where $\operatorname{\mathbf{Sel}}^{1}(Y)$ denotes the subfamily of absolute integrable selections; see [2, 1, 7].
Definition (1.1) sets the expectation of a random set to be the set of all expectations of its selections. Similar concepts for the median of a random set or for event probabilities can be defined as collections of these operators for all possible selections of the random set (see Subsection 1.3). In this paper, we study restricted selection sets — subsets of $\operatorname{\mathbf{Sel}}(Y)$ defined by imposing scalar constraints on the selections. Concretely, we consider sets of the form
$\operatorname{\mathbf{Sel}}(Y\mid\kappa):=\{y\in\operatorname{\mathbf{Sel}}(Y):\mathbb{E}[y]=\kappa\},\qquad\operatorname{\mathbf{Sel}}(Y\mid m):=\{y\in\operatorname{\mathbf{Sel}}(Y):m\in\operatorname*{Med}(y)\}.$
and more generally, selection sets restricted by moment or quantile conditions. Our aim is to understand (i) when such restricted selection sets are nonempty, (ii) how the constraint narrows the set of selections relative to $\operatorname{\mathbf{Sel}}(Y)$, and (iii) which values of basic scalar functionals,
$\mathbb{E}[y],\qquad med(y),\qquad\mathbf{P}(y\in A),$
can be attained by selections $y$ satisfying a given constraint.
We develop the core theory in the one-dimensional setting, where $Y=[y_{L},y_{U}]$ is a random interval with $y_{L}\leq y_{U}$ $\mathbf{P}$-a.s. In this case the Aumann expectation has the simple representation $\mathbb{E}[Y]=[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})]$, so that for every $y\in\operatorname{\mathbf{Sel}}^{1}(Y)$,
$\mathbb{E}(y_{L})\leq\mathbb{E}[y]\leq\mathbb{E}(y_{U}).$ (1.2)
Similarly, since each $y\in\operatorname{\mathbf{Sel}}(Y)$ stochastically dominates $y_{L}$ and is stochastically dominated by $y_{U}$,
$med(y_{L})\leq med(y)\leq med(y_{U}).$ (1.3)
The central question is: how do the attainable ranges in (1.2) and (1.3) change when one restricts attention to a nonempty subset $\mathcal{S}\subset\operatorname{\mathbf{Sel}}(Y)$ defined by a scalar constraint?
The non-atomicity of $(\Omega,\mathcal{A},\mathbf{P})$ allows us to use measurable selection and rearrangement arguments in a sharp way. At a technical level, several of our proofs combine classical properties of the Aumann expectation with Hardy–Littlewood type inequalities and quantile representations. This yields explicit and, in many cases, closed-form descriptions of the extremal values of expectations, medians, event probabilities, and higher moments over constrained selection sets.
We then show that the one-dimensional theory is sufficient for the general case: the selection set of any random compact convex set in $\mathbb{R}^{d}$ can be approximated in $L^{1}$ by selection sets of disjoint unions of random cubes, each of which decomposes coordinate-wise into a product of one-dimensional interval selection sets. This reduces every constrained selection problem for random convex sets in $\mathbb{R}^{d}$ to the interval case.
@@ 1.1. Related Literature
Our work is rooted in the classical theory of random closed sets and selection expectation. The notion of the Aumann integral for set-valued mappings was introduced in [2] and developed further in [1], where convexity and support-function representations play a central role. In the one-dimensional setting considered here, the random sets are intervals $Y=[y_{L},y_{U}]$, and the identity $\mathbb{E}[Y]=[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})]$ is a special case of these general results. For the basic theory of random closed sets, measurable selections, and the measurable graph theorem, we refer to [5, 3, 7].
Rearrangement and quantile methods appear throughout analysis and probability in the study of extremal inequalities. Our arguments for quantile attainability and median-restricted expectations are in the spirit of classical Hardy-Littlewood-Pólya inequalities, as presented for instance in [6, 4]. We employ these tools in a simple random interval setting, where they yield explicit formulas for the effect of median constraints on the expectation and for extremal event probabilities under a mean restriction.
Random sets and Aumann expectations have been used in a variety of applied contexts, including mathematical economics, finance, and statistics; see, for example, the monograph [7] and the references therein. In particular, models with set-valued or interval-valued observations naturally lead to constrained selection problems of the type studied here. The present paper is primarily structural: we develop the core results in a one-dimensional probability-theoretic setting and then show how they extend to random compact convex sets in $\mathbb{R}^{d}$ via a dyadic approximation of the selection set.
@@ 1.2. Our Contribution
We develop a detailed theory of restricted selection sets for random closed sets on a non-atomic probability space. Starting from the benchmark Aumann expectation range and median interval for the full selection set $\operatorname{\mathbf{Sel}}(Y)$, we first show that these bounds are sharp: every point of $\mathbb{E}[Y]=[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})]$ is realized as $\mathbb{E}[y]$ for some $y\in\operatorname{\mathbf{Sel}}(Y)$ (Proposition 2.2). More generally, for each $\alpha\in(0,1)$, any value between the capacity and containment quantiles of $Y$ can be attained as the $\alpha$–quantile of a selection (Proposition 2.4).
We then analyze how scalar constraints narrow the selection set and sharpen these bounds. In the median-restricted case, we obtain sharp bounds on the attainable means: under mild feasibility conditions, the set
@@
We also examine the converse: does a mean constraint $\mathbb{E}[y]=\kappa$ narrow the set of attainable medians? A two-state construction (Example 2.11) shows that the answer is negative in general: every point in the benchmark median interval
$\operatorname*{med}_{-}(y_{L}),\,\operatorname*{med}_{+}(y_{U})]$
can remain attainable as a median under the mean constraint. For event probabilities $\mathbf{P}(y\in A)$, we refine the classical random-set bounds in Proposition 2.12 by imposing a mean constraint and deriving a dual (Lagrangian) representation (Theorem 2.13). This leads to an explicit threshold selection rule (Theorem 2.14) and a closed-form formula for the extremal probabilities $\mathbf{P}(y\in A)$ over $\operatorname{\mathbf{Sel}}(Y\mid\kappa)$ (Corollary 2.15), while preserving consistency with the unrestricted capacity/containment bounds (Proposition 2.16).
We further extend the framework in three directions. First, we consider higher-moment constraints and obtain dual representations for the extremal means under an $r$th-moment restriction (Theorem 2.20). Second, we treat quantile constraints at general levels $\alpha\neq\tfrac{1}{2}$ and show that the induced set of attainable means is always a convex subset of the Aumann range (Proposition 2.23). Finally, we show that the selection set of any random compact convex set $Y\subset\mathbb{R}^{d}$ can be approximated in $L^{1}$ by selection sets of disjoint unions of random cubes, giving us an approximation-based reduction of constrained selection problems to finite families of one-dimensional interval-selection problems (Theorem 2.34).
@@ 1.3. Notation and preliminaries
Spaces and sets. We denote by $\mathbb{R}$ the real line and, for $d\in\mathbb{N}$, by $\mathbb{R}^{d}$ the dd–dimensional Euclidean space equipped with its usual topology and Euclidean norm $\|\cdot\|$. Let $\mathcal{K}(\mathbb{R})$ be the family of all nonempty closed subsets of $\mathbb{R}$. Our random sets always take values in $\mathcal{K}(\mathbb{R})$. For general background on random closed sets we refer to [7].
@@
where $\mathcal{B}(\mathcal{K}(\mathbb{R}))$ is the Borel $\sigma$–algebra generated by the Fell (hit–or–miss) topology (see [7]). In this note we consider the special case of bounded random intervals, that is,
@@
The existence of measurable selections for closed-valued measurable multifunctions is classical, see e.g. [5, 3]. We denote the family of all measurable selections of $Y$ by
@@
Following [2], the Aumann expectation (selection expectation) of a random closed set $Y$ with at least one integrable selection is defined as the closed set
@@
where the closure is taken in the usual topology on $\mathbb{R}$; see also [1, 7]. For a random interval $Y=[y_{L},y_{U}]$ satisfying (1.4), one has the simple representation
@@
see [7]. In the one–dimensional setting we will mainly use the half–line sets $(-\infty,t]$, $t\in\mathbb{R}$, and write
@@
We write $\operatorname*{Med}(z)$ for the set of all such medians. In the scalar case, $\operatorname*{Med}(z)$ is always a nonempty closed interval (possibly a singleton). We also use
$\operatorname*{med}_{-}(z):=\inf\operatorname*{Med}(z),\qquad\operatorname*{med}_{+}(z):=\sup\operatorname*{Med}(z)$
for the lower and upper endpoints of the median set.
@@
2. Main Results
In this section, we study restricted selection sets of a random interval $Y=[y_{L},y_{U}]$ on the non-atomic probability space $(\Omega,\mathcal{A},\mathbf{P})$ introduced in Subsection 1.3. First, we record benchmark ranges for the expectation, median, and event probabilities of measurable selections of $Y$. Then we investigate how these ranges are sharpened by simple scalar restrictions on the mean or median of the selection. Throughout we assume $y_{L},y_{U}$ are measurable real-valued random variables with $\mathbf{P}(y_{L}\leq y_{U})=1$ and
@@ 2. Main Results
2.1. Benchmark bounds
Let $y\in\operatorname{\mathbf{Sel}}(Y)$ be an arbitrary measurable selection of the random interval $Y=[y_{L},y_{U}]$. The following bounds for the expectation and median of a selection $y$ are immediate.
@@ 2.1. Benchmark bounds
$\mathbb{E}[y]\in\Theta_{E}:=\big[\mathbb{E}(y_{L}),\mathbb{E}(y_{U})\big].$ (2.1)
Second, since $y_{L}\leq y\leq y_{U}$ almost surely, every median of $y$ must lie between the lower endpoint of the median set of $y_{L}$ and the upper endpoint of the median set of $y_{U}$. Thus
$\operatorname*{Med}(y)\subseteq\Theta_{M}:=\big[\operatorname*{med}_{-}(y_{L}),\,\operatorname*{med}_{+}(y_{U})\big].$ (2.2)
where the right-hand side is understood as the smallest interval containing both $\operatorname*{Med}(y_{L})$ and $\operatorname*{Med}(y_{U})$.
@@ 2.2. Restricted selection sets
Let $Y=[y_{L},y_{U}]$ be a random interval with measurable endpoints satisfying $y_{L}\leq y_{U}$ a.s. and assume $y_{L},y_{U}\in L^{1}(\mathbf{P})$, the space of integrable random variables. For each $t\in[0,1]$ define
@@ 2.3. Quantile attainability via selections
Let $(\Omega,\mathcal{F},\mathbf{P})$ be a non-atomic probability space and let $Y$ be a (measurable) random closed convex set in $\mathbb{R}$ (a random interval). For $t\in\mathbb{R}$, define
@@
On $A^{-}\cup A^{0}$, the set $Y^{\leq m}(\omega)$ is nonempty; on $A^{0}\cup A^{+}$, the set $Y^{>m}(\omega)$ is nonempty (note that on $A^{0}$ we have $y_{U}(\omega)>m$, since $y_{U}(\omega)\leq m$ together with $y_{L}(\omega)\leq m$ would place $\omega$ in $A^{-}$). Because $Y$ has a measurable graph and the half-lines are Borel, $\operatorname{graph}(Y^{\leq m})$ and $\operatorname{graph}(Y^{>m})$ are measurable. By standard measurable selection results (e.g. via a Castaing representation, see [5, 3]) we can take measurable selections
@@
Since $Y(\omega)=[y_{L}(\omega),y_{U}(\omega)]$ is an interval, we have the following explicit forms:
$z^{-}(\omega)=\begin{cases}y_{U}(\omega),&\omega\in A^{-},\\[2.0pt] m,&\omega\in A^{0},\end{cases}$
(noting that $\sup Y^{\leq m}(\omega)=m$ on $A^{0}$ and $y_{U}(\omega)$ on $A^{-}$), while on $A^{0}\cup A^{+}$, every selection $z^{+}$ of $Y^{>m}$ satisfies
$z^{+}(\omega)>m.$ (2.3)
On $A^{+}$ we take $z^{+}(\omega)=y_{L}(\omega)>m$. On $A^{0}$ we set $z^{+}(\omega)\in(m,y_{U}(\omega)]$, so in particular $z^{+}(\omega)>m$.
For notational convenience, we complete $z^{-}$ and $z^{+}$ to selections of $Y$ by setting
@@
Taking $\alpha=\tfrac{1}{2}$, the previous proposition implies that every point in
$\operatorname*{med}_{-}(y_{L}),\,\operatorname*{med}_{+}(y_{U})]$
is a median of some selection of $Y$, so the benchmark median interval $\Theta_{M}$ is sharp.